# InversionView: A General-Purpose Method for Reading Information from Neural Activations

### v2

@@ Abstract
The inner workings of neural networks can be better understood if we can fully decipher the information encoded in neural activations. In this paper, we argue that this information is embodied by the subset of inputs that give rise to similar activations. Computing such subsets is nontrivial as the input space is exponentially large. We propose InversionView, which allows us to practically inspect this subset by sampling from a trained decoder model conditioned on activations. This helps uncover the information content of activation vectors, and facilitates understanding of the algorithms implemented by transformer models. We present three case studies where we investigate models ranging from small transformers to GPT-2. In these studies, we demonstrate the characteristics of our method, show the distinctive advantages it offers, and provide causally verified circuits.
@@ 1 Introduction
We showcase the usefulness of the method in three case studies: a character counting task, Indirect Object Identification, and 3-digit addition. The character counting task illustrates how the method uncovers how information is processed and forgotten in a small transformer. In Indirect Object Identification in GPT2-Small [39], we use InversionView to easily interpret the information encoded in the components identified by [39], substantially simplifying the interpretability workflow. For 3-digit addition, we use InversionView to provide for the first time a fully verified circuit sufficient for performing the task. Across the case studies, InversionView allows us to rapidly generate hypotheses about the information encoded in each activation site. Coupled with attention patterns or patching methods, we reverse-engineer the flow of information, which we verify using causal interventions.
@@ 2.2 Conditional Decoder Model
Directly enumerating $B_{\mathbf{z}^{q},f,\epsilon}$ is in general not scalable, as the input space grows exponentially with the sequence length. To efficiently inspect $B_{\mathbf{z}^{q},f,\epsilon}$, we train a conditional decoder model that takes as input the activation $\mathbf{z}^{q}$ and generates inputs giving rise to similar activations in the model under investigation. In the following, we refer to the original model that we are interpreting as the probed model, the conditional decoder as the decoder, the place in the probed model from which we take the activation as the activation site (e.g., the output of $i$th layer), the inputs generated by the decoder as samples, and the index of a token in the sequence as position.
We implement the decoder as an autoregressive language model conditioned on $\mathbf{z}^{q}$, decoding input samples ${\bf x}$. As the decoder’s training objective corresponds to recovering ${\bf x}$ exactly, sampling at temperature 1 will typically not cover the full $\epsilon$-preimage. We increase diversity by drawing samples at higher temperatures and with noise added to $\mathbf{z}^{q}$ (details in Appendix A.2). We then evaluate $D(f({\bf x}),{\bf z}^{q})$ at each position in each sample ${\bf x}$, select the position minimizing $D$,111 In some cases, including Figure 2(b) 3(b), we fix a particular position for interpretation. See Appendix A.3. determine membership in $B_{\mathbf{z}^{q},f,\epsilon}$, and subsample in-$\epsilon$-preimage and out-of-$\epsilon$-preimage samples for inspection.
An important question is whether this method, relying on a black-box decoder, produces valid $\epsilon$-preimages. As we evaluate $D(f({\bf x}),{\bf z}^{q})$ for each sample ${\bf x}$, correctness (are all generated samples in the $\epsilon$-preimage?) is ensured by design. We explicitly verify completeness (are all elements of the $\epsilon$-preimage generated at reasonable probability?) by enumerating inputs in one of our case studies (Appendix B). Another, more scalable, approach is to design counter-examples ${\bf x}$ not satisfying a hypothesis about the content of $B_{\mathbf{z}^{q},f,\epsilon}$, and checking if they in fact are outside of it. In our experiments, we found that these examples were always outside of $B_{\mathbf{z}^{q},f,\epsilon}$.
@@ 3.1 Setup and Background
Using notation based on [14] and [29], we denote the residual stream [14] as $x^{i,\{\mathrm{pre},\mathrm{mid},\mathrm{post}\}}\in\mathbb{R}^{N\times d}$, where $i$ is the layer (an attention (sub)layer + an MLP (sub)layer) index, $N$ is the number of input tokens, $d$ is the model dimension, $\mathrm{pre}$, $\mathrm{mid}$, $\mathrm{post}$ stand for the residual stream before the attention layer, between attention and MLP layer, and after the MLP layer. For example, $x^{0,\mathrm{pre}}$ is the sum of token and position embedding, $x^{0,\mathrm{mid}}$ is the sum of the output of the first attention layer and $x^{0,\mathrm{pre}}$, and $x^{0,\mathrm{post}}$ is the sum of the output of the first MLP layer and $x^{0,\mathrm{mid}}$. Note that $x^{i,\mathrm{post}}=x^{i+1,\mathrm{pre}}$. We use subscript $t$ to refer to the activation at token position $t$, e.g., $x^{i,\mathrm{mid}}_{t}\in\mathbb{R}^{d}$. The attention layer output decomposes into outputs of individual heads $h^{i,j}(\cdot)$, i.e., $x^{i,\mathrm{mid}}=x^{i,\mathrm{pre}}+\sum_{j}h^{i,j}(\mbox{LN}(x^{i,\mathrm{pre}}))$, where $\mbox{LN}(\cdot)$ represents layer normalization (GPT style/pre-layer-norm). We denote the attention head’s output as $a^{i,j}$, i.e., $a^{i,j}=h^{i,j}(\mbox{LN}(x^{i,\mathrm{pre}}))$.
@@
We train a single two-layer transformer decoder across all activation sites of interest. The query activation ${\bf z}^{q}$ is concatenated with an activation site embedding ${\bf e}$, a learned embedding layer indicating where the activation comes from, passed through multiple MLP layers with residual connections, and then made available to the attention heads in each layer of the decoder, alongside the already present tokens from the input, so that each attention head can also attend to the post-processed query activation in addition to the context tokens. Appendix C has technical details. Each training example is a triple consisting of an activation vector $\in\mathbb{R}^{d}$, the activation site index, and the input, on which the decoder is trained with a language modeling objective.
@@ 3.2 Character Counting
We causally verified our hypothesis using activation patching [38, 16] on (position, head output) pairs. As the attention head in layer 1 attends almost entirely to the target character, only heads $a^{0,0}_{tc}$, $a^{0,0}_{:}$, and $a^{1,0}_{:}$ can possibly play a role in routing count information. We patch their outputs with activations from a contrast example flipping a single character before “|”. To verify dependencies, we patch cumulatively, starting either at the lowest or highest layer, with some fixed ordering within each layer. By the end of patching, the model prediction is on the contrast example by the end of patching. When patching a head, we attribute to it the increment in the difference of $LD$ before and after patching, where $LD$ denotes the logit difference between original count and the count in the contrast example. Cumulative patching allows us to verify that all paths are accounted for; unaccounted paths through a layer would manifest in an increment at a point where all hypothesized connected activations in the other layer have already been patched. Results (Figure 3(a)) show that patching either of the heads in the hypothesized path ($a^{0,0}_{tc}$ and $a^{1,0}_{:}$) is sufficient to absorb the entire effect on logit differences, confirming the hypothesis. See Appendix D.3 for further details and D.4 for further experiments.
@@ 3.3 IOI circuit in GPT-2 small
To test the applicability of InversionView to transformers pretrained on real-world data, we apply our method to the activations in the indirect object identification (IOI) circuit in GPT-2 small [34] discovered by Wang et al. [39]. We apply InversionView to the components of the circuit, read out the information, and compare it with the information or function that Wang et al. [39] had ingeniously inferred using a variety of tailored methods, such as patching and investigating effects on logits and attention. We show that InversionView unveils the information contained in the activation sites, with results agreeing with those of Wang et al. [39], while avoiding the need for tailored methods.
@@
We start with the Name Mover Head 9.9, which Wang et al. [39] found moves the IO name to the residual stream of END. 3(b) shows the $\epsilon$-preimage at “to”. The samples in the $\epsilon$-preimage share the name “Justin” as the IO. The head also shows similar activity at some other positions (Appendix A.3). Results are consistent across query inputs. Therefore, the result of InversionView is consistent with the conclusions of Wang et al. [39] on head 9.9. Applying the same analysis to other heads (Table 2), we recovered information in high agreement with the information that [39] had ingeniously inferred using a variety of tailored methods. Among the 26 attention heads they identified, InversionView indicates a different interpretation in only 3 cases; these (0.1, 0.10, 5.9) were challenging for the methods used before (Appendix E.3). In summary, InversionView scales to larger models.
@@
We next applied InversionView to the problem of adding 3-digit numbers, between 100 and 999. Input strings have the form “B362+405=767E” or “B824+692=1516E”, and are tokenized at the character level. We use F1, F2, F3 to denote the three digits of the first operand and S1, S2, S3 for the digits of the second operand, and A1, A2, A3, A4 (if it exists) for the three or four digits of the answer, and C2, C3 for the carry from tens place and ones place (i.e., C2: whether F2+S2$\geq$10, C3: whether F3+S3$\geq$10). Unlike Quirke and Barez [33], we do not pad numbers to have all the same length; hence, positional information is insufficient to determine the place value of each digit.
@@
Figure 6: 3-Digit Addition Task: InversionView uncovers different ways in which digit representation is encoded in activations. Left: The digits in the hundreds place are encoded separately and hence generations denote them as separate entities. Right: The digits in the tens place are encoded as a sum (9 in this case) and the generations represent different 2-partitions (7+2, 6+3, 1+8, 5+4, etc.) of that sum.
@@ 4 Discussion
Traditional probing [e.g. 2, 6] trains a supervised classifier to assess how much information about a variable of interest is encoded in an activation site. It requires a hypothesis in advance and is thus inherently limited to hypotheses conceived a priori by the researcher. InversionView, on the other hand, helps researchers form hypotheses without any need for prior guesses, and allows fine-grained per-activation interpretation. Inspecting attention patterns [e.g. 10] is a traditional approach to inferring information flow, and we have drawn on it in our analyses. More recently, path patching [39, 20] causally identifies paths along which information flows. In contrast to these approaches, InversionView aims at aiding generation of hypotheses about what information is contained in activations and passed along. While the information flow provides an upper bound on the information passed along by tracing back to the input token, it is insufficient for determining how information is processed and abstracted. For instance, in Section 3.2, occurrences of the target character are causally connected to $a^{0,0}_{tc}$, which then connects to $a^{1,0}_{:}$ (direct or mediated by MLP layer 0). Without looking at encoded information, we only know that the information in these paths is related to the occurrences of the target character, but not whether it is their identity, positions, count, etc. More generally, when a component reads a component that itself has read from multiple components, connectivity does not tell us which pieces of information are passed on. We provide further discussion in Appendix H.
@@
InversionView offers distinctive advantages and makes analyses feasible that are otherwise very hard to do with other methods. It can also improve the interpretability workflow in coordination with other methods. For example, one may first use methods such as path patching or attribution [15] to localize activity to specific components, and then understand the function of these components using InversionView. In sum, InversionView is worth adding to the toolbox of interpretability research.
@@
InversionView relies on a black-box decoder, which needs to be trained using relevant inputs and whose completeness needs to be validated by counter-examples. Also, InversionView, while easing the human’s task, is still not automated, and interpretation can be laborious when there are many activation sites. We focus on models up to 125M parameters; scaling the technique to models with billions of parameters is an interesting problem for future work, which will likely require advances in localizing behavior to a tractable number of components of interest. Fourth, interpretation uses a metric $D(\cdot,\cdot)$. The geometry, however, in general could be nonisotropic and more complex than the $L^{2}$-based metrics we used. We leave the exploration of this to future work.
4.1 Related Work
Trained probing classifiers are arguably the most common method for uncovering information from activations [e.g. 2, 5, 4, 40, 37, 24, 25]. More recent methods obtain insights by projecting representations into the vocabulary space [30, 7, 31, 22, 18], patching them into an LLM [19], or by passing them through sparse feature vectors [9, 36, 12]. Dar et al. [13] directly interprets model parameters by projecting them into the embedding space. Another important area of research is uncovering information flow. Geva et al. [18] intervene on attention weights to study information flow. Activation patching, a causal intervention method, can be used to study the causal effect of an activation on the output, and can help localize where information is stored [28, 35], or find alignment between a high-level causal model and inner states of a neural model [16, 17, 41]. While activation patching affects all downstream computation in the model, path patching [39, 20] restricts effects for downstream components in order to identify the path along which information flows [39, 11, 21, 26]. Recent work has started using LLMs to generate interpretations [8, 9].

### v3

@@ Abstract
The inner workings of neural networks can be better understood if we can fully decipher the information encoded in neural activations. In this paper, we argue that this information is embodied by the subset of inputs that give rise to similar activations. Computing such subsets is nontrivial as the input space is exponentially large. We propose InversionView, which allows us to practically inspect this subset by sampling from a trained decoder model conditioned on activations. This helps uncover the information content of activation vectors, and facilitates understanding of the algorithms implemented by transformer models. We present four case studies where we investigate models ranging from small transformers to GPT-2. In these studies, we demonstrate the characteristics of our method, show the distinctive advantages it offers, and provide causally verified circuits. [^1] Code is available at https://github.com/huangxt39/InversionView
@@ 1 Introduction
We showcase the usefulness of the method in three thorough case studies: a character counting task, Indirect Object Identification, and 3-digit addition. We also present preliminary results on the factual recall task, demonstrating the applicability of our method to larger models. The character counting task illustrates how the method uncovers how information is processed and forgotten in a small transformer. In Indirect Object Identification in GPT2-Small [50], we use InversionView to easily interpret the information encoded in the components identified by [50], substantially simplifying the interpretability workflow. For 3-digit addition, we use InversionView to provide for the first time a fully verified circuit sufficient for performing the task. Across the case studies, InversionView allows us to rapidly generate hypotheses about the information encoded in each activation site. Coupled with attention patterns or patching methods, we reverse-engineer the flow of information, which we verify using causal interventions.
@@ 2.2 Conditional Decoder Model
In this paper, we study the setting where $\mathbf{x}^{q}$ is a sequence. Directly enumerating $B_{\mathbf{z}^{q},f,\epsilon}$ is in general not scalable, as the input space grows exponentially with the sequence length. To efficiently inspect $B_{\mathbf{z}^{q},f,\epsilon}$, we train a conditional decoder model that takes as input the activation $\mathbf{z}^{q}$ and generates inputs giving rise to similar activations in the model under investigation. In the following, we refer to the original model that we are interpreting as the probed model, the conditional decoder as the decoder, the place in the probed model from which we take the activation as the activation site (e.g., the output of $i$th layer), the inputs generated by the decoder as samples, and the index of a token in the sequence as position.
We implement the decoder as an autoregressive language model conditioned on $\mathbf{z}^{q}$, decoding input samples ${\bf x}$. As the decoder’s training objective corresponds to recovering ${\bf x}$ exactly, sampling at temperature 1 will typically not cover the full $\epsilon$-preimage. Thus, for generating elements of the $\epsilon$-preimage, we increase diversity by drawing samples at higher temperatures and with noise added to $\mathbf{z}^{q}$ (details in Appendix A.2). Note that the activation $f({\bf x})$ will depend on the position in the input sequence; hence, we separately evaluate $D(f({\bf x}),{\bf z}^{q})$ at each position $i$ in each sample ${\bf x}$, and select the position $i$ minimizing $D(f({\bf x}),{\bf z}^{q})$,333 In some cases, including Figure 2(b) 3(b), we fix a particular position for interpretation. See Appendix A.3. determine membership in $B_{\mathbf{z}^{q},f,\epsilon}$, and subsample in-$\epsilon$-preimage and out-of-$\epsilon$-preimage samples for inspection.
An important question is whether this method, relying on a black-box decoder, produces valid $\epsilon$-preimages. We can approach this question from two perspectives, namely correctness (are all generated samples in the $\epsilon$-preimage?) and completeness (are the samples representative of the $\epsilon$-preimage?). As we evaluate $D(f({\bf x}),{\bf z}^{q})$ for each generated samples, correctness is ensured by design, thus we avoid making conclusion based on wrong samples. As for completeness, we ask whether all elements of the $\epsilon$-preimage can be output by the method at reasonable probability. If some groups of inputs in $\epsilon$-preimage are systematically missing from the generations, one may overestimate the information contained in activations. But this behavior would be punished by the training objective, since the loss on these examples would be high. We explicitly verify completeness by enumerating inputs in one of our case studies (Appendix B). Another approach is to design counter-examples ${\bf x}$ not satisfying a hypothesis about the content of $B_{\mathbf{z}^{q},f,\epsilon}$, and checking if they in fact are outside of it. In our experiments, we found that these examples were always outside of $B_{\mathbf{z}^{q},f,\epsilon}$.
@@ 3.1 Setup and Background
In the transformer architecture, outputs from each layer are added to their inputs due to residual connection. The representations of each token are only updated by additive updates, forming a residual stream [17]. Using notation based on [17] and [39], we denote the residual stream as $x^{i,\{\mathrm{pre},\mathrm{mid},\mathrm{post}\}}\in\mathbb{R}^{N\times d}$, where $i$ is the layer (an attention (sub)layer + an MLP (sub)layer) index, $N$ is the number of input tokens, $d$ is the model dimension, $\mathrm{pre}$, $\mathrm{mid}$, $\mathrm{post}$ stand for the residual stream before the attention layer, between attention and MLP layer, and after the MLP layer. For example, $x^{0,\mathrm{pre}}$ is the sum of token and position embedding, $x^{0,\mathrm{mid}}$ is the sum of the output of the first attention layer and $x^{0,\mathrm{pre}}$, and $x^{0,\mathrm{post}}$ is the sum of the output of the first MLP layer and $x^{0,\mathrm{mid}}$. Note that $x^{i,\mathrm{post}}=x^{i+1,\mathrm{pre}}$. We use subscript $t$ to refer to the activation at token position $t$, e.g., $x^{i,\mathrm{mid}}_{t}\in\mathbb{R}^{d}$. The attention layer output decomposes into outputs of individual heads $h^{i,j}(\cdot)$, i.e., $x^{i,\mathrm{mid}}=x^{i,\mathrm{pre}}+\sum_{j}h^{i,j}(\mbox{LN}(x^{i,\mathrm{pre}}))$, where $\mbox{LN}(\cdot)$ represents layer normalization (GPT style/pre-layer-norm). We denote the attention head’s output as $a^{i,j}$, i.e., $a^{i,j}=h^{i,j}(\mbox{LN}(x^{i,\mathrm{pre}}))$.
@@
We train a single two-layer transformer decoder across all activation sites of interest; we chose this architecture for simplicity. The query activation ${\bf z}^{q}$ is concatenated with an activation site embedding ${\bf e}$, a learned embedding layer indicating where the activation comes from, passed through multiple MLP layers with residual connections, and then made available to the attention heads in each layer of the decoder, alongside the already present tokens from the input, so that each attention head can also attend to the post-processed query activation in addition to the context tokens. Each training example is a triple consisting of an activation vector ${\bf z}^{q}\in\mathbb{R}^{d}$, the activation site index, and the input, on which the decoder is trained with a language modeling objective. Appendix C has technical details.
@@ 3.2 Character Counting
We causally verified our hypothesis using activation patching [49, 21] on (position, head output) pairs. As the attention head in layer 1 attends almost entirely to the target character, only head outputs $a^{0,0}_{tc}$, $a^{0,0}_{:}$, and $a^{1,0}_{:}$ can possibly play a role in routing count information. We patch their outputs with activations from a contrast example flipping a single character before “|”. We patch activations cumulatively, starting either at the lowest or highest layer, with some fixed ordering within each layer. For example, we patch $a^{0,0}_{:}$ and observe how final logits change compared to the clean run, then we patch both $a^{0,0}_{:}$ and $a^{0,0}_{tc}$ and do the same, and so forth. By the end of patching, the model prediction will be flipped. When adding an activation to the patched set, we attribute to it the increment in the difference of $LD$ before and after patching, where $LD$ denotes the logit difference between original count and the count in the contrast example. Cumulative patching allows us to observe dependencies: For instance, as we hypothesize that $a^{1,0}_{:}$ is completely dependent on $a^{0,0}_{tc}$, we expect that, when $a^{0,0}_{tc}$ is already patched, patching $a^{1,0}_{:}$ will have no further effect, whereas when $a^{0,0}_{tc}$ is not patched, patching $a^{1,0}_{:}$ will have a significant effect. Results (Figure 3(a)) match our prediction: Patching either of the activation in the hypothesized path ($a^{0,0}_{tc}$ and $a^{1,0}_{:}$) is sufficient to absorb the entire effect on logit differences, confirming the hypothesis. See Appendix D.3 for further details and D.4 for further experiments.
@@ 3.3 IOI circuit in GPT-2 small
To test the applicability of InversionView to transformers pretrained on real-world data, we apply our method to the activations in the indirect object identification (IOI) circuit in GPT-2 small [44] discovered by Wang et al. [50]. We apply InversionView to the components of the circuit, read out the information, and compare it with the information or function that Wang et al. [50] had ingeniously inferred using a variety of tailored methods, such as patching and investigating effects on logits and attention. We show that InversionView unveils the information contained in the attention heads’ outputs, with results agreeing with those of Wang et al. [50].
@@
We start with the Name Mover Head 9.9, which Wang et al. [50] found moves the IO name to the residual stream of END. 3(b) shows the $\epsilon$-preimage at “to”. The samples in the $\epsilon$-preimage share the name “Justin” as the IO. The head also shows similar activity at some other positions (Appendix A.3). Results are consistent across query inputs. Therefore, the result of InversionView is consistent with the conclusions of Wang et al. [50] on head 9.9. Applying the same analysis to other heads (Table 2), we recovered information in high agreement with the information that [50] had inferred using multiple tailored methods. For example, Wang et al. [50] found S-Inhibition heads were outputting both token signals (value of S) and position signals (position of S1) by patching these heads’ outputs from a series of counterfactual datasets. These datasets are designed to disentangle the two effects, in which token and/or position information are ablated or inverted. On the other hand, these two kinds of information can be directly read out by InversionView (Figure 18 shows an example for one of the S-Inhibition heads that contains position information), and there is no need to guess the possible information to design patching experiments. Overall, among the 26 attention heads that Wang et al. [50] identified, InversionView indicates a different interpretation in only 3 cases; these (0.1, 0.10, 5.9) were challenging for the methods used before (Appendix E.3). In summary, InversionView scales to larger models.
@@
We next applied InversionView to the problem of adding 3-digit numbers, between 100 and 999. Input strings have the form “B362+405=767E” or “B824+692=1516E”, and are tokenized at the character level. We use F1, F2, F3 to denote the three digits of the first operand and S1, S2, S3 for the digits of the second operand, and A1, A2, A3, A4 (if it exists) for the three or four digits of the answer, and C2, C3 for the carry from tens place and ones place (i.e., C2: whether F2+S2$\geq$10, C3: whether F3+S3$\geq$10). Unlike [43], we do not left-pad answers to have all the same length; hence, positional information is insufficient to determine the place value of each digit.
@@
Figure 6: 3-Digit Addition Task: InversionView uncovers different ways in which digit is encoded in activations. Left: The digits in the hundreds place are encoded separately and hence generations denote them as separate entities. Right: The digits in the tens place are encoded as a sum (9 in this case) and the generations represent different 2-partitions (7+2, 6+3, 1+8, 5+4, etc.) of that sum.
@@
3.5 Factual Recall
To test whether InversionView can be applied to larger language models, we explore how GPT-2 XL (1.5B parameters) performs the task of recalling factual associations. In this case study, our intention is not to provide a full interpretation of the computations performed to solve this task, which we deem out of scope for this paper. Instead, we show that InversionView produces interpretable results on larger models by focusing on a relatively small set of important attention heads in upper layers. The decoder model in this case study is based on GPT-2 Medium, because we expect a more complex inverse mapping from activation to inputs to be learned. We observe the resulting $\epsilon$-preimage can express high-level knowledge (Figure 38-43), and sometimes can predict the failure of the model (Appendix H.6). Using InversionView, we again shed light on the underlying mechanism of the model. We present detailed findings in Appendix H.
@@ 4 Discussion
Traditional probing [e.g. 2, 6] trains a supervised classifier to assess how much information about a variable of interest is encoded in an activation site. It requires a hypothesis in advance and is thus inherently limited to hypotheses conceived a priori by the researcher. InversionView, on the other hand, helps researchers form hypotheses without any need for prior guesses, and allows fine-grained per-activation interpretation. Inspecting attention patterns [e.g. 12] is a traditional approach to inferring information flow, and we have drawn on it in our analyses. More recently, path patching [50, 26] causally identifies paths along which information flows. In contrast to these approaches, InversionView aims at aiding generation of hypotheses about what information is contained in activations and passed along. While the information flow provides an upper bound on the information passed along by tracing back to the input token, it is insufficient for determining how information is processed and abstracted. For instance, in Section 3.2, occurrences of the target character are causally connected to $a^{0,0}_{tc}$, which then connects to $a^{1,0}_{:}$ (direct or mediated by MLP layer 0). Without looking at encoded information, we only know that the information in these paths is related to the occurrences of the target character, but not whether it is their identity, positions, count, etc. More generally, when a component reads a component that itself has read from multiple components, connectivity does not tell us which pieces of information are passed on. In addition, Direct Logit Attribution (DLA) [50, 23, 40] is also frequently used for inspecting the output information of transformer’s components. We argue that DLA is only suitable for studying model components that directly affect model’s final output. For those components whose effect is mediated by other components, their output information is meant to be read by a downstream component, thus not necessarily visible when projecting to the vocabulary space. We provide further discussion in Appendix I.
@@
InversionView offers distinctive advantages and makes analyses feasible that are otherwise very hard to do with other methods. It can also improve the interpretability workflow in coordination with other methods. For example, one may first use methods such as path patching or attribution [46, 18] to localize activity to specific components, and then understand the function of these components using InversionView. In sum, InversionView is worth adding to the toolbox of interpretability research.
@@
InversionView relies on a black-box decoder, which needs to be trained using relevant inputs and whose completeness needs to be validated by counter-examples. Also, InversionView, while easing the human’s task, is still not automated, and interpretation can be laborious when there are many activation sites. We focus on models up to 1.5B parameters; scaling the technique to large models is an interesting problem for future work, which will likely require advances in localizing behavior to a tractable number of components of interest. Fourth, interpretation uses a metric $D(\cdot,\cdot)$. The geometry, however, in general could be nonisotropic and treating each dimension equally could be sub-optimal. We leave the exploration of this to future work.
5 Related Work
Trained probing classifiers are arguably the most common method for uncovering information from activations [e.g. 2, 5, 4, 51, 48, 32, 33]. More recent methods obtain insights by projecting representations into the vocabulary space [40, 7, 41, 30, 24], patching them into an LLM [25], or by decomposing them into interpretable sparse feature vectors [10, 47, 14]. Dar et al. [16] directly interprets model parameters by projecting them into the embedding space. Another important area of research is uncovering information flow. Geva et al. [24] intervene on attention weights to study information flow. Activation patching, a causal intervention method, can be used to study the causal effect of an activation on the output, and can help localize where information is stored [36, 45], or find alignment between a high-level causal model and inner states of a neural model [21, 22, 53]. While activation patching affects all downstream computation in the model, path patching [50, 26] restricts effects for downstream components in order to identify the path along which information flows [50, 13, 27, 34]. Recent work has started using LLMs to generate interpretations [8, 10].